# Do machines fail like humans? A human-centred out-of-distribution spectrum for mapping error alignment

### 08 Mar 2026 v1

@@ 2.3.1 Patterns and variability in model–human alignment across OOD levels
To examine how alignment changes over distortion severity and OOD levels, we plot and describe accuracy, EC, and MA patterns across the full range of distortion levels for three distortion types (Figure 4, High-Pass, Low-Pass and Contrast; see Figure S9 and S10 for all other distortion types). In general, we observe high variability across OOD levels which suggests analyzing OOD regimes as separate groups.
Accuracy performance and alignment decline as we move from the reference distribution through near- to far-OOD – models broadly track the human trend but differ in where and how they deviate. In the reference regime (yellow), models typically exceed human accuracy, though Contrast does not have a condition that falls within this region. For some curves, we see an increase in EC in the near OOD region compared to the reference distribution. It is unclear precisely why this is and further investigation may be required. MA is also less informative here due to very few errors. Entering near-OOD (green) regime, behavior begins to separate by distortion type. For High-Pass, some families keep pace with or exceed human accuracy while others do not. Humans generally outperform most models for Low-Pass, whereas most models have higher accuracy than humans for Contrast. Across distortion types, human–human EC remains higher than human–model EC. Importantly, EC slightly increases despite reduced accuracy for High-Pass and Contrast, indicating that people fail on a shared set of samples as perceptual difficulty increases within this regime. MA provides a complementary view: across all distortion types, several models approach human–human MA, meaning that their incorrect predictions resemble human errors. Notably, MA is a stricter criterion than EC, so numerical differences tend to be smaller than in EC.
In the far-OOD regime (blue), under Low-Pass, humans often outperform models, whereas High-Pass and Contrast largely preserve their near-OOD ordering. Human–human EC under High-Pass becomes more variable (shaded error bars) yet mostly remains above human–model EC. For Low-Pass, human–human EC stays higher, though its margin over human–model EC can narrow. Of note, EC for humans for Contrast is exceptionally high (see level C05), reflecting strong agreement about some images even at higher distortion levels. MA under Contrast is comparatively low, meaning that both humans and models may agree on the correct-incorrect boundary (high EC) but diverge on their mistakes. By contrast, Low-Pass produces the largest gap between human–human and model–human MA.
Although alignment degrades with OOD level, the degree of degradation is strongly model-dependent. VLMs tend to maintain higher MA, ViTs struggle under Low-Pass, and ViT-based architectures like Swin and MaxViT show robustness under High-Pass. However, these patterns vary across OOD regimes, making it difficult to draw conclusions about alignment across the full spectrum. Therefore, we focus on quantifying regime-specific alignment in Section 2.3.2 and 2.3.3.
@@ 2.3.2 Architectural family effects on human alignment
Based on the qualitative observation on Figure 4 (see also Figure S9, S10), models belonging to the same architectural family tend to exhibit characteristic error-alignment profiles. This motivates a natural question: are models architecturally closer to one another, such as those within the same sub-family (e.g., VGGs, ResNets) or within the same broader super-family (e.g., CNNs, ViTs), also closer in how they align with humans? We formalise two hypotheses: H1) model-human alignment is more similar within sub-families than within super-families; H2) model-human alignment is more similar within super-families than across super-families.
We test these hypotheses using permutation tests on model-human alignment profiles. For each model, we concatenate EC and MA values across all distortion types into a single alignment vector and compute pairwise Euclidean distances between models. For H1, we compare within-sub-family distances (e.g., VGG-16 vs. VGG-19) to across-sub-family distances within the same super-family (e.g., VGGs vs. ResNets). For H2, we compare within-super-family distances to across-super-family distances. The p-values and the effect size are reported in Table 2. At the super-family level, under both OOD regimes, this effect is statistically robust ($p<0.01$) and with strong effect sizes, suggesting that broad design principles reliably influence model–human alignment. At the sub-family level, results depends upon architectures and OOD regimes.
In general, far-OOD leads to a stronger effect, indicating greater alignment similarity within a model family. For VLMs, intra-family similarity is significantly higher than inter-family similarity. CNNs strongly reject the null hypothesis in both near- and far-OOD regimes, showing the strongest effect sizes among the three superfamilies. By contrast, ViTs do not exhibit statistically reliable sub-family clustering in either regime, suggesting that architectural variations within the ViT family may give rise to more diverse error patterns.
@@ 2.3.3 Profiles of model-human error alignment over distortion types
To further examine how architectural families influence model–human alignment across distortion types, we select one representative condition for each distortion type from both near- and far-OOD regimes and visualise their alignment profiles using radar plots (Figure 5). Similarity in the shape of the resulting curves indicates similar error patterns between systems, enabling intuitive comparison across architectures. We focus here on results at the model superfamily level; results for individual subfamilies are shown in Figure S12.
In the near-OOD regime, the human–human EC curve is roughly circular with relatively narrow error bars, indicating similar behavioral agreement across distortion types. In general, VLMs and CNNs present higher EC than ViTs. ViTs only surpass CNNs under Low-Pass distortions. For MA, human–human alignment is less stable, showing wider error bars and greater variability across distortion types. Several distortion types yield similar EC but markedly different MA patterns. For instance, Uniform Noise and Low-Pass produce comparable and high human-human EC, but the former yields much higher MA. Conversely, EidolonII produces comparatively low EC but exceptionally high MA for both human–human and human–model pairs, implying the same incorrect decisions are often made for the co-errors. CNNs achieve both high EC and MA under Uniform Noise. ViTs, by contrast, exhibit both low EC and MA under Uniform Noise. VLMs display particularly strong human alignment (especially MA) under High-Pass and Low-Pass, albeit with substantial variability. All three superfamilies show similarly strong MA under Power-Equalisation.
In the far-OOD regime, human–human EC becomes less uniform across distortion types. The EC for Contrast and Uniform Noise changes only slightly from near- to far-OOD. In contrast, High-Pass and Low-Pass produce the largest declines in human–human EC. Among model families, CNNs generally exhibit the lowest EC across distortions, except under Contrast, where they remain competitive. Under Low-Pass and Uniform Noise, CNNs’ EC values approach zero. ViTs show a distinct advantage in the three Eidolon conditions, whereas VLMs achieve exceptionally high EC under Low-Pass. For MA, human–human alignment decreases considerably compared with the near-OOD regime, suggesting that under these highly impaired conditions, it becomes much harder for humans to exhibit shared misperceptions. Nonetheless, Low-Pass, Stylized, and EidolonIII maintain comparatively high human–human MA, while Contrast leads to minimal agreement on incorrect labels. Across models, VLMs achieve particularly high MA under Phase-Scrambling and Stylized. ViTs still attain MA values close to those of VLMs in several conditions, although the absolute values remain low. CNNs consistently show very low MA across all distortions.
@@ 2.3.4 Model-human alignment rankings shifts over OOD regimes reflect different inductive biases
Finally, we examine which models showed the most human-like error patterns by ranking average model-human alignment scores. To compare how models align with humans across different model families, we normalise model–human alignment by human–human alignment. Specifically, for each model $m$ and condition $c$, we compute the alignment ratio
$\rho_{m}^{c}=\frac{a_{m}^{c}}{a^{c}}$ (2)
where $a^{c}_{m}$ is the model’s average alignment with humans (mean value of EC and MA; see Figure S11 for EC and MA separately) and $\hat{a}^{c}$ is the corresponding human–human alignment score for that condition. This ratio accounts for the inherent variability in human agreement under each distortion and evaluate models relative to a human baseline. We note averaging $\rho_{m}^{c}$ over all representative conditions within an OOD regime as $\bar{\rho}_{m}$, a summary statistic for ranking models in that regime. The highest-ranked model in a regime is therefore the one whose alignment with humans is most comparable to the agreement observed among humans themselves.
The rankings of $\bar{\rho}_{m}$ for the models in near- and far-OOD regimes are shown in Figure 6A and the alignment ratios for each model in near and far-OOD regime are presented in Figure 6B. We find considerable differences in model-human alignment between near- and far-OOD: many models that align closely with humans in near-OOD no longer rank highly in far-OOD, showing that perceptual difficulty reshapes the human-based behavioral consistency of models (Figure 6A). In the near-OOD regime, VLMs and CNNs tend to achieve a higher alignment ratio than ViTs ($p^{\prime}s<0.01$), although ViTs typically outperform CNNs in standard recognition tasks.
In the far-OOD regime, VLMs maintain high alignment rankings, but to a similar level to ViTs, both of which had higher human alignment over CNNs ($p<0.01$; Figure 6C). If VLMs exhibit greater model-human alignment due to language and semantic knowledge from their text-based training, this suggests that semantic information is marked affected and even completely removed in far-OOD regimes.
@@ 3 Discussion
Model-human alignment comparisons have typically treated different distortion types and levels similarly, despite being distinct in nature. Our human-centered framework addresses this issue by defining an OOD spectrum that groups conditions by human perceptual difficulty, rather than by arbitrary parameter-based severity levels. We show that model-human alignment depends on the type and level of distortion, where vision-language models are the most consistently human-aligned across both near- and far-OOD conditions, but CNNs are more aligned than ViTs for near OOD and ViTs are more aligned than CNNs for far OOD.
@@ 3.1 Why a Human-Centred OOD Spectrum
When models are exposed to distributional shifts, the way their performance degrades can reveal their inductive biases – the visual features they exploit for recognition. The error profile across distortion types thus serves as a behavioral signature of a system’s information-processing strategies. However, drawing valid conclusions requires a principled way to compare distortions. Distortion parameters are incommensurable across types, and prior approaches, including aggregating across severity levels, qualitatively selecting comparable distortions, or normalising to a reference model  [17, 6, 24], each introduce limitations. Our perceptual OOD spectrum grounded in human perceptual difficulty addresses this by quantifying and grouping conditions by deviation from human performance under standard, baseline performance, rather than deviation from training statistics or arbitrary parameters. The resulting OOD spectrum provides a principled basis for assessing whether models fail like humans across the full spectrum of perceptual challenge.
First, this framework allows meaningful comparisons at matched levels of human perceptual difficulty across distortion types. Without such matching, apparent differences in alignment may simply reflect task difficulty rather than genuine architectural biases. For example, Eidolon I parameter 8, High-Pass parameter 1, and Low-Pass parameter 5 induce similar levels of behavioral shift from baseline, even though they appear visually very different. Second, it enables the identification of appropriate conditions for evaluating model-human alignment. Applying this method reveals four regimes: reference, representing natural variation on undistorted or lightly distorted images; near-OOD, representing moderate accuracy reduction; far-OOD, a transitional zone where performance declines at varying rates across distortion types; and extreme-OOD, where images lack sufficient information for recognition and performance falls to or below chance. Qualitatively speaking, reference, near-OOD, and extreme-OOD emerged as strongly coherent components, while far-OOD is more diffuse, likely due to sparse sampling in this range. When both humans and models fail entirely, alignment values become uninformative. We therefore exclude extreme-OOD for evaluation. Prior work that aggregated extreme-OOD conditions may have led to conclusions that are not well supported under finer-grained OOD analysis.
@@ 3.2 Establishing the Human–Human Behavioral Alignment Baseline
Human-human alignment patterns across the OOD spectrum strongly suggest that near-OOD and far-OOD regimes should be evaluated separately, as systems likely rely on different strategies across these conditions. In the near-OOD regime, humans show roughly uniform EC across distortions, suggesting that errors may primarily target "harder" undistorted images – those that become ambiguous under minor degradation. Such errors depend more on the stimulus than the observer. Human–human MA is also relatively high for near-OOD, meaning distortions which obscure the true class are also likely to suggest a specific alternative, a "false lead" that consistently draws multiple observers toward the same incorrect response, reflecting strong stimulus-driven effects.
@@
In the far-OOD regime, this pattern shifts. EC degrades gracefully, indicating that individual differences in responses increases as distortions become more severe, reflecting observer dependence rather than purely stimulus effects. MA degrades more quickly than EC, suggesting that while false leads are still present, they become less common. This suggests that stronger distortions disrupt recognition in different ways from system to system. At extreme-OOD, both metrics flatline, implying no meaningful class information remains. MA in particular indicates that false leads are almost entirely absent. The absence of both accurate predictions and systematic error biases reflects a lack of meaningful human–human alignment in this regime where human–model alignment cannot be meaningfully assessed. Researchers should therefore avoid these conditions when drawing conclusions.
@@
These regime-specific patterns can be understood in terms of how EC and MA relate to stimulus space. EC and MA inform us about the alignment between two systems’ decision boundaries in stimulus space – the degree to which two systems agree, but they are sensitive to the data distribution on which they are computed [11, 44]. As different distortion levels produce distinct data distributions occupying different regions of stimulus space, computing EC and MA across pooled distortion levels would conflate alignment patterns that may differ substantially across these regions. Our findings support this concern: EC and MA vary considerably across distortion levels within the same distortion type, demonstrating that pooling across severity levels would obscure meaningful variation in alignment.
EC and MA reveal how humans align within a given distortion condition, but not whether error patterns are similar across conditions. CLED addresses this by comparing behavioral responses across non-overlapping stimulus sets. The implication is that distortion conditions with low CLED impair the visual features to which humans are sensitive in similar ways. This can arise by chance due to the sparsity of errors when both conditions are easy (high accuracy). If they are both hard and CLED is low, the relevant visual features upon which humans depend to make decisions may have been impaired in similar ways. If one condition is easy and another hard, then the hard distortion would need to somehow balance the impairment evenly across the visual features used for classification to produce a similar error pattern. This seems less plausible, and we would therefore expect distortions of different levels of difficulty to typically have higher CLEDs than pairs of similar difficulty. Both the t-SNE visualisation and quantitative analysis confirm this: OOD level exerts a stronger effect on human error structure than distortion type.
@@
Taken together, these findings suggest that humans employ different decision-making strategies at different levels of perceptual difficulty: in near-OOD, errors are stimulus-driven and consistent across observers, while in far-OOD, errors become more observer-dependent and less systematic. To explore whether models align with humans in their information-processing strategies, we must therefore evaluate these regimes separately – a model that aligns with human errors in near-OOD may not do so in far-OOD, and vice versa.

### 11 Apr 2026 v2

@@ 2.3.1 Patterns and variability in model–human alignment across OOD levels
To examine how alignment changes over distortion severity and OOD levels, we plot accuracy, EC, and MA patterns across the full range of distortion levels for three distortion types (Figure 3, High-Pass, Low-Pass and Contrast; see Figure S8 and S9 for other distortion types). In general, we observe high variability across OOD levels.
Accuracy performance and alignment decline as we move from the reference distribution through near- to far-OOD – models broadly track the human trend but differ in where and how they deviate. In the reference regime (yellow background), models typically exceed human accuracy, though Contrast does not have a condition that falls within this region. For some curves, we see an increase in EC in the near OOD region compared to the reference distribution. It is unclear precisely why this is and further investigation may be required. MA is also less informative here due to there being very few errors. In the near-OOD regime (green background) human–human EC remains higher than model-human EC across distortion types. Importantly, EC slightly increases despite reduced accuracy for High-Pass and Contrast, indicating that people fail on a shared set of samples as perceptual difficulty increases within this regime. MA provides a complementary view: across all distortion types, several models approach human–human MA, meaning that their incorrect predictions resemble human errors. Notably, MA is a stricter criterion than EC, so numerical differences tend to be smaller than in EC.
Entering far-OOD regime (blue), under Low-Pass, humans often outperform models, whereas High-Pass and Contrast largely preserve their near-OOD ordering. Human–human EC under High-Pass becomes more variable (shaded error bars) yet mostly remains above model-human EC. For Low-Pass, human–human EC stays higher, though its margin over model-human EC narrows. Of note, EC for humans for Contrast is exceptionally high (see level C05), reflecting strong agreement about some images even at higher distortion levels. MA under Contrast is comparatively low, meaning that both humans and models may agree on the correct-incorrect boundary (high EC) but diverge on their mistakes. By contrast, Low-Pass produces the largest gap between human–human and model–human MA.
Although alignment degrades with OOD level, the degree of degradation is strongly model-dependent. VLMs tend to maintain higher MA, CNNs struggle under Low-Pass, and ViT-based architectures like Swin and MaxViT show robustness under High-Pass. However, these patterns vary across OOD regimes, making it difficult to draw conclusions about alignment across the full spectrum. Therefore, we focus on quantifying regime-specific alignment in the following sections.
@@ 2.3.2 Architectural family effects on human alignment
Based on qualitative observation of Figure 3 (see also Figure S 8, S 9), models belonging to the same architectural family tend to exhibit characteristic error-alignment profiles. This motivates a natural question: are models architecturally closer to one another also more similar in how they align with humans? For example, are the two models from the VGG family more similar in their alignment than they are to other members of the wider CNNs? We formalise two hypotheses: H1) model-human alignment is more similar between members of the same sub-family than between members of the same super-family; H2) model-human alignment is more similar between members of the same super-family than across super-families. Here, we define two levels of architectural families: sub-family (e.g., VGGs, ResNets) and super-family (e.g., CNNs, ViTs).
We test these hypotheses using permutation tests on model-human alignment profiles. For each model, we concatenate EC and MA values across all distortion types into a single alignment vector and compute pairwise Euclidean distances between models. For H1, we compare within-sub-family distances (e.g., VGG-16 vs. VGG-19) to across-sub-family distances within the same super-family (e.g., VGGs vs. ResNets). For H2, we compare within-super-family distances to across-super-family distances. The $p$-values and the effect size are reported in Table 2. At the super-family level, under both OOD regimes, this effect is statistically robust ($p<0.01$) and with strong effect sizes. At the sub-family level, results depends upon architectures and OOD regimes. Generally, far-OOD leads to a stronger effect, indicating greater alignment similarity within a model family. CNNs strongly reject the null hypothesis in both near- and far-OOD regimes, showing the strongest effect sizes among the three superfamilies. By contrast, ViTs do not exhibit statistically reliable sub-family clustering in either regime, suggesting that architectural variations within the ViT family may give rise to more diverse error patterns. For VLMs, the effect in far-OOD is higher than that in near-OOD.
@@ 2.3.3 Profiles of model-human error alignment over distortion types
To further examine how architectural families influence model–human alignment across distortion types, we select one representative condition for each distortion type from both near- and far-OOD regimes and visualize their alignment profiles using radar plots (Figure 4). Similarity in the shape of the resulting curves indicates similar error patterns between architectures across distortion types. We focus here on results at the model superfamily level; results for individual subfamilies are shown in Figure S11.
In the near-OOD regime, the human–human EC curve is roughly circular with relatively narrow error bars, indicating similar behavioral agreement across distortion types. In general, VLMs and CNNs present higher EC than ViTs. ViTs only surpass CNNs under Low-Pass distortions. For MA, human–human alignment is less stable, showing wider error bars and greater variability across distortion types. Several distortion types yield similar EC but markedly different MA patterns. For instance, Uniform Noise and Low-Pass produce comparable and high human-human EC, but the former yields much higher MA. Conversely, EidolonII produces comparatively low EC but exceptionally high MA for both human–human and model-human pairs. CNNs achieve both high EC and MA under Uniform Noise, where ViTs do not. VLMs display particularly strong human alignment (especially MA) under High-Pass and Low-Pass, albeit with substantial variability. All three superfamilies show similarly strong MA under Power-Equalisation.
In the far-OOD regime, human–human EC becomes less uniform across distortion types. The EC for Contrast and Uniform Noise changes only slightly from near- to far-OOD. In contrast, High-Pass and Low-Pass produce the largest declines in human–human EC. Among model families, CNNs generally exhibit the lowest EC across distortions, except under Contrast, where they remain competitive. Under Low-Pass and Uniform Noise, CNNs’ EC values approach zero. ViTs show a distinct advantage in the three Eidolon conditions, whereas VLMs achieve exceptionally high EC under Low-Pass. For MA, human–human alignment decreases considerably compared with the near-OOD regime, suggesting that under these highly impaired conditions, it becomes much harder for humans to exhibit shared misclassifications. Nonetheless, Low-Pass, Stylized, and EidolonIII maintain comparatively high human–human MA, while Contrast leads to minimal agreement on incorrect labels. Across models, VLMs achieve particularly high MA under Phase-Scrambling and Stylized. ViTs still attain MA values close to those of VLMs in several conditions, although the absolute values remain low. CNNs consistently show very low MA across all distortions.
@@ 2.3.4 Model-human alignment rankings shifts over OOD regimes reflect different inductive biases
Finally, we examine which models showed the most human-like error patterns by ranking average model-human alignment scores. To compare how models align with humans across different model families, we normalise model–human alignment measures by human–human alignment on the corresponding condition. Specifically, for each model $m$ and condition $c$, we compute the alignment ratio
$\rho_{m}^{c}=\frac{a_{m}^{c}}{\hat{a}^{c}}$ (2)
where $a^{c}_{m}$ is the model’s average alignment with humans (mean value of EC and MA; see Figure S5 for EC and MA separately) and $\hat{a}^{c}$ is the corresponding human–human alignment score for that condition. This ratio accounts for the inherent variability in human agreement under each distortion and evaluate models relative to a human baseline. We denote the average of $\rho_{m}^{c}$ over all representative conditions within an OOD regime as $\bar{\rho}_{m}$, a summary statistic for ranking models in that regime. The highest-ranked model in a regime is therefore the one whose alignment with humans is most comparable to the agreement observed among humans themselves.
The rankings of $\bar{\rho}_{m}$ for the models in near- and far-OOD regimes are shown in Figure 5A and the alignment ratios for each model in near and far-OOD regime are presented in Figure 5B. We find considerable differences in model-human alignment between near- and far-OOD: many models that align closely with humans in near-OOD no longer rank highly in far-OOD, showing that perceptual difficulty reshapes the human-based behavioral consistency of models (Figure 5A). In the near-OOD regime, VLMs and CNNs tend to achieve a higher alignment ratio than ViTs ($p^{\prime}s<0.01$), although ViTs typically outperform CNNs in terms of accuracy in the reference conditions. In the far-OOD regime, VLMs maintain high alignment rankings, but to a similar level to ViTs, both of which had higher human alignment over CNNs ($p<0.01$; Figure 5C).
@@ 3 Discussion
Model-human alignment comparisons have typically treated different distortion types and levels similarly, despite being distinct in nature. Our human-centred framework addresses this by defining an OOD spectrum that groups conditions by human perceptual difficulty, rather than by arbitrary severity levels. We show that model-human alignment depends on the type and level of distortion. VLMs are the most consistently human-aligned across both near- and far-OOD conditions, whereas CNNs are more aligned than ViTs for near OOD and ViTs are comparable to VLMs and more aligned than CNNs for far OOD.
@@ 3.1 Why a Human-Centred OOD Spectrum
When models are exposed to distributional shifts, the way their performance degrades can reveal their inductive biases – the visual features they exploit for recognition. The error profile across distortion types thus serves as a behavioral signature of a system’s information-processing strategies. However, drawing valid conclusions requires a principled way to compare distortions. Distortion parameters are incommensurable across types, and prior approaches, including aggregating across severity levels, qualitatively selecting comparable distortions, or normalising to a reference model  [17, 6, 24], each have their limitations. We addresses this by quantifying and grouping conditions by deviation from human performance under standard, baseline performance. The resulting OOD spectrum provides a principled basis for assessing whether models fail like humans across the full spectrum of perceptual challenge.
First, this framework allows meaningful comparisons at matched levels of human perceptual difficulty across distortion types. Without this, apparent differences in alignment may simply reflect task difficulty rather than genuine architectural biases. For example, Eidolon I parameter 8, High-Pass parameter 1, and Low-Pass parameter 5 induce similar levels of behavioral shift from baseline, even though they appear visually very different. Second, it enables the identification of appropriate conditions for evaluating model-human alignment. The method reveals four regimes: reference, representing natural variation on undistorted or lightly distorted images; near-OOD, representing moderate accuracy reduction; far-OOD, a transitional zone where performance declines at varying rates across distortion types; and extreme-OOD, where images lack sufficient information for recognition and performance falls to or below chance. Qualitatively, reference, near-OOD, and extreme-OOD emerged as strongly coherent components, while far-OOD is more diffuse, likely due to sparse sampling in this range. When both humans and models fail entirely, alignment values become uninformative. We therefore exclude extreme-OOD for evaluation. Prior work that aggregated extreme-OOD conditions may have led to conclusions that are not well supported under finer-grained OOD analysis.
@@ 3.2 Establishing the Human–Human Behavioral Alignment Baseline
Human-human alignment patterns across the OOD spectrum strongly suggest that near-OOD and far-OOD regimes should be evaluated separately, as systems likely rely on different strategies across these conditions. In the near-OOD regime, humans show roughly uniform EC across distortions, suggesting that errors may primarily target “harder" undistorted images – those that become ambiguous under minor degradation. Such errors depend more on the stimulus than the observer. Human–human MA is also relatively high for near-OOD, meaning distortions which obscure the true class are also relatively more likely to suggest a specific alternative, a “false lead" that consistently draws multiple observers toward the same incorrect response, reflecting strong stimulus-driven effects.
@@
In the far-OOD regime, this pattern shifts. EC degrades gracefully, indicating that individual differences in responses increase as distortions become more severe, reflecting observer dependence rather than stimulus-driven effects. MA degrades more quickly than EC, suggesting that while false leads are still present, they become less common. This suggests that stronger distortions disrupt recognition in different ways across systems. At extreme-OOD, both metrics flatline, implying no meaningful class information remains. MA in particular indicates that false leads are almost entirely absent. The absence of accurate predictions and systematic errors reflects a lack of meaningful human–human alignment where model-human alignment cannot be meaningfully assessed. Researchers should therefore avoid these conditions when drawing conclusions about model-human alignment.
@@
These regime-specific patterns can be understood in terms of how EC and MA relate to stimulus space. EC and MA inform us about the alignment between two systems’ decision boundaries in stimulus space. They measure the degree to which two systems agree on where the distinctions between classes are, but are sensitive to where in stimulus space the concentration of stimuli sit [11, 44]. Different distortion levels, and to a lesser degree different types, produce distinct data distributions and occupy different regions of stimulus space. Computing EC and MA across pooled conditions would conflate alignment patterns that may differ substantially across these regions.
CLED patterns further support the importance of perceptual difficulty, though the effect varies across regimes. In the near-OOD regime, conditions at similar difficulty levels produce similar error profiles regardless of distortion type, suggesting that perceptual difficulty dominates over the specific features impaired, although some of this may arise by chance due to the sparsity of errors when both conditions have high accuracy. In the far-OOD regime, some conditions within the same distortion type tend to cluster together, indicating that distortion-specific feature impairment may play a larger role when degradation is severe. When two conditions are both difficult and CLED is low, the relevant visual features upon which humans depend for recognition may have been impaired in similar ways.
@@
Taken together, these findings suggest that humans employ different decision-making strategies at different levels of perceptual difficulty: in near-OOD, errors are stimulus-driven and more consistent across observers, while in far-OOD, errors become more observer-dependent and less systematic. To explore whether models align with humans in their information-processing strategies, we must therefore evaluate these regimes separately – a model that aligns with human errors in near-OOD may not do so in far-OOD, and vice versa.